OpenStudy (anonymous):

Which statement describes the translation of y= –5(x2 + 2) – 3 from standard position? Moved up and to the right. Moved up and to the left. Moved down and to the left. Moved down and to the right.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

C or D

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

it is? :D

hero (hero):

yummydum you know this stuff already. Besides, even if you didn't, all you'd have to do is graph this function and the parent function, then compare both.

OpenStudy (anonymous):

What is the range of the graph of y = 3(x – 3)2 + 4? y >= 4 y <= 4 y >= 3 y <= 3

hero (hero):

Have you tried graphing the function yet? That would be the easiest way to figure it out

OpenStudy (anonymous):

y=3x^2-18x+31

hero (hero):

You didn't need to do that to graph it

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

yes or no???

OpenStudy (anonymous):

@Chlorophyll

hero (hero):

You got it, but you should be a little more sure of yourself when trying to figure out the answer.

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

Of course you're correct, because 3(x – 3)2 >=0 => 4 is minimum Thus the range y ≥ 4